OpenStudy (anonymous):

X1 belongs to (0, 1) xn+1=xn^3-3xn^2+3xn where n + 1, n are indices * Study the monotony. As detailed as pissible please.

OpenStudy (anonymous):

Since \(x_1\) belongs to (presumably) the interval \((0,1)\), then \(x_1\) is some positive rational number between 0 and 1. This means you have \[x_2={x_1}^3-3{x_1}^2+3x_1\] To describe the monotony of a recursive sequence, you must compare successive terms. For the particular recursion, you want to be able to compare the effect of higher powers of the \(n\)-th term. For example, suppose \(x_1=\dfrac{1}{2}\). Higher powers of this paticular \(x_1\) will be smaller than the first power, i.e. \({x_1}^2=\dfrac{1}{4}\) and \({x_1}^3=\dfrac{1}{8}\), so \(x_1>{x_1}^2>{x_1}^3>\cdots\). This is true for any number \(x_1\) in \((0,1)\). Keeping with the example of \(x_1=\dfrac{1}{2}\), we get \[x_2=\frac{1}{8}-\frac{3}{4}+\frac{3}{2}=\frac{7}{8}>\frac{1}{2}\] You should be able to see that the \(3x_n\) term will always overcome the size of the \({x_n}^3-3{x_n}^2\) terms, so the sequence is increasing (and actually caps at 1).